LASER SPECTROSCOPY OF NEODYMIUM MONXIDE.

\underline{C. LINTON} , Physics Department, University of New Brunswick, Fredericton, NB, Canada E3B 5A3; C. EFFANTIN, P. CROZET, A. J. ROSS, J. D′INCAN, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment A. Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France; E. A. SHENYAVSKAYA, Institute for High Energy Density of Associated Institute for High Temperatures of Russian Academy of Science, Izorskaya St. 13/19, Moscow 125412, Russia.

Laser induced fluorescence spectra of ^{142}NdO have been excited using both fixed frequency argon ion and tunable ring dye lasers and detected at high resolution with a Fourier transform spectrometer. Nine low lying electronic states resulting from the Nd^{+2}(4f^{3}6s)O^{−2} configuration were detected of which four, the second lowest Ω = 2, 3 and 5 states, (2)2, (2)3, (2)5, and the lowest Ω = 6 state, (1)6, have been observed for the first time. In addition, new vibrational levels were observed in the lowest Ω = 5, (1)5 (v = 1) and second lowest Ω = 4, (2)4 (v = 1, 2) states. Abnormally large doubling in both Ω = 2 states has been attributed to interactions involving neighboring Ω = 1 and 0 states. Several perturbations were observed and used as an aid in assigning some of the states. Both the order and energies of the low lying states have been shown to be consistent with Ligand Field theory calculations. Rotational relaxation in several of the spectra has allowed calculation of accurate rotational constants for several states while, for other states, approximate parameters have been calculated from combination differences.